OpenStudy (anonymous):

Please help What is the sum of the arithmetic sequence 134, 122, 110 …, if there are 32 terms?

OpenStudy (tpaulus):

Find the general form, then use that to find the last term. then use the sum of an arithmetic sequence equation to find the sum.\[s = n/2 * (n1 + n32)\]

OpenStudy (campbell_st):

well the sum is simply \[S_{n} = \frac{n}{2}[2a + (n -1)d]\] a = 1st term, n = number of terms and d = common difference in your question you know n = 32, the 1st term is 134 find the common difference. ..d and then substitute and evaluate. the value of d will be negative

OpenStudy (anonymous):

–1,664 –1,632 –1,600 –1,568

OpenStudy (anonymous):

Those are the answers.^

OpenStudy (anonymous):

S = [134 - (0)(12)] + [134 - (1)(12)] + [134 - (2)(12)] + . . . + [134 - (31)(12)] S = (134)(32) - (12)(1 + 2 + 3 + . . . + 31) S = (134)(32) - (12)(31)(32)/2

OpenStudy (anonymous):

All good now, @maria-17 ?

OpenStudy (anonymous):

THanksss.

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @maria-17

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

Sure, why not? Best though to close this one and just start another post.